OpenStudy (anonymous):

Using complete sentences, explain how to graph x + 3y = 3 using the x- and y-intercepts method.

OpenStudy (anonymous):

let x=0, solve for y let y=0, solver for x now you have two points that determine line... just draw it passing through those two points.

OpenStudy (anonymous):

Maybe this will help? x + 3y = 3 0 + 3y = 3 y = 1 x + 0 = 3 x = 3 Then just graph it.